# Mean-field theory of 1+1D $\mathbb{Z}_{2}$ lattice gauge theory with matter

### v1

@@ A.2 Including the $\mathbb{Z}_{2}$ electric field term $h$
$\hat{\tau}^{x}_{j,j+1}=e^{i\pi\sum_{l<j}\hat{n}_{l}}.$ (27)
The above expression is obtained by considering open boundary conditions where we assume that the chain starts and ends with an anti-string ($\tau^{x}=+1$). By fixing the Gauss law to the physical sector ($\hat{G}_{j}\ket{\psi}=+1\ket{\psi}$) we can express $\mathbb{Z}_{2}$ electric field on site $j$ as a product of Gauss law operators
@@ A.3 Mapping between the Kitaev chain and the transverse-field Ising model
^$\hat{\mathcal{H}}_{I}=-J\sum_{j}\hat{\sigma}^{x}_{j}\hat{\sigma}^{x}_{j+1}-h_{z}\sum_{j}\hat{\sigma}^{z}_{j},$ (29)
@@
$\hat{\sigma}^{z}_{j}\rightarrow\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1},\quad\hat{\sigma}^{x}_{j}\hat{\sigma}^{x}_{j+1}\rightarrow\hat{\tau}^{z}_{j,j+1},$ (30)
@@
^$\hat{\mathcal{H}}_{I}=-J\sum_{j}\hat{\tau}^{z}_{j,j+1}-h_{z}\sum_{j}\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}.$ (31)
@@ Appendix B Mapping of the $\mathbb{Z}_{2}$ LGT to the spin-1/2 model
$\hat{n}_{j}=\frac{1}{2}\left(1-\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}\right).$ (33)
@@
$\begin{split}\hat{a}^{\dagger}_{j}&=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\frac{1}{2}\left(1+\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}\right),\\ \hat{a}_{j}&=\Bigg{(}\prod_{l<j}\hat{\tau}^{z}_{l,l+1}\Bigg{)}\frac{1}{2}\left(1-\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}\right).\end{split}$ (34)
We once again note that we always consider that our chain starts and ends with a link. The logic of the operator is following: we first take into account that the particle at some site $j$ can be added only if the site $j$ is empty, otherwise the state has to be annihilated. This is achieved by first applying the operator 12($\frac{1}{2}\left(1+\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}\right)$ which projects to the state with an empty site $j$. After this we apply the product of $\prod_{l<j}\hat{\tau}^{z}_{l,l+1}$ operators up to the link which attaches to our site from the left. In such way we create a domain wall and thus a particle at that site. The same logic applies for the annihilation operator.
@@
^$\hat{\mathcal{H}}_{LGT}^{s}=t\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}-\hat{S}^{z}_{j}\right)-h\sum_{j=1}^{L}2\hat{S}^{x}_{j},$ (35)
@@
^$\hat{\mathcal{H}}_{\lambda}^{s}=\lambda\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}+\hat{S}^{z}_{j}\right)$ (36)
and a term which is proportional to the chemical potential ^$\hat{\mathcal{H}}_{\mu}^{s}=\mu\sum_{j=1}^{L-1}2\hat{S}^{x}_{j}\hat{S}^{x}_{j+1}$. This term is important for the DMRG simulations, since it is used to control the filling in the lattice.
@@ C.1 Derivation of the mean-field models
^$\hat{\mathcal{H}}=-t\sum_{j=1}^{L-1}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}_{j}\rangle+\langle\hat{c}^{\dagger}_{j}\hat{c}_{j+1}\rangle\right)\hat{\tau}^{z}_{\langle j,j+1\rangle}+\lambda\sum_{j=1}^{L-1}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}\rangle+\langle\hat{c}_{j}\hat{c}_{j+1}\rangle\right)\hat{\tau}^{z}_{\langle j,j+1\rangle}-h\sum_{j=0}^{L}\hat{\tau}^{x}_{\langle j,j+1\rangle}$
[…]
@@
where we defined $t_{c}=t\langle\hat{\tau}^{z}_{\langle j,j+1}\rangle$ and $\lambda_{c}=\lambda\langle\hat{\tau}^{z}_{\langle j,j+1}\rangle$, and added the chemical potential term $\mu_{c}$ which enforces the desired filling of the chain.
@@
^$\hat{\mathcal{H}}_{\tau}=-g\sum_{j=1}^{L-1}\hat{\tau}^{z}_{\langle j,j+1\rangle}-h\sum_{j=0}^{L}\hat{\tau}^{x}_{\langle j,j+1\rangle}$
$+\mu_{\tau}\sum_{j=1}^{L}\left(\hat{\tau}^{x}_{j-1,j}\hat{\tau}^{x}_{j,j+1}-\left(1-2n\right)\right),$ (42)
@@ C.2 Solving the mean-field theory for the charges
^$\hat{\mathcal{H}}_{c}=\sum_{k>0}\left(\mu_{c}-2t_{c}\cos(k)\right)\left(\hat{c}^{\dagger}_{k}\hat{c}_{k}+\hat{c}^{\dagger}_{-k}\hat{c}_{-k}\right)$ (45)
@@
^$\hat{\mathcal{H}}_{c}=\sum_{k>0}\begin{pmatrix}\hat{c}^{\dagger}_{k}&\hat{c}_{-k}\end{pmatrix}\begin{pmatrix}\epsilon(k)&-2i\lambda_{c}\sin(k)\\ 2i\lambda_{c}\sin(k)&-\epsilon(k)\end{pmatrix}\begin{pmatrix}\hat{c}_{k}\\ \hat{c}^{\dagger}_{-k}\end{pmatrix}$
@@
^$\hat{\mathcal{H}}_{c}-\sum_{k>0}\epsilon(k)+\mu_{c}nL=\sum_{k>0}\begin{pmatrix}\hat{c}^{\dagger}_{k}&\hat{c}_{-k}\end{pmatrix}\bm{H}(k)\begin{pmatrix}\hat{c}_{k}\\ \hat{c}^{\dagger}_{-k}\end{pmatrix}\\ =\sum_{k>0}\begin{pmatrix}\hat{b}^{\dagger}_{k}&\hat{b}_{-k}\end{pmatrix}\bm{\Lambda}(k)\begin{pmatrix}\hat{b}_{k}\\ \hat{b}^{\dagger}_{-k}\end{pmatrix},$ (49)
@@
^$\hat{\mathcal{H}}_{c}=\sum_{k}\Lambda_{+}\hat{b}^{\dagger}_{k}\hat{b}_{k}-\sum_{k>0}\Lambda_{+}+\sum_{k>0}\epsilon(k)-\mu_{c}nL.$ (51)
@@ C.3 Solving the mean-field theory for the gauge fields
$E_{0}=-t_{c}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}_{j}\rangle+\langle\hat{c}^{\dagger}_{j}\hat{c}_{j+1}\rangle\right)\\ +\lambda_{c}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}\rangle+\langle\hat{c}_{j}\hat{c}_{j+1}\rangle\right)=-\langle\hat{\tau}^{z}_{\langle j,j+1}\rangle g.$ (56)
@@ D.1 DMRG simulation of the $\mathbb{Z}_{2}$ LGT
^$\hat{\mathcal{H}}_{\rm DMRG}=t\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}-\hat{S}^{z}_{j}\right)\\ +\lambda\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}+\hat{S}^{z}_{j}\right)\\ -h\sum_{j=1}^{L}2\hat{S}^{x}_{j}+\mu\sum_{j=1}^{L-1}2\hat{S}^{x}_{j}\hat{S}^{x}_{j+1},$ (59)
@@
Figure 14: Entanglement entropy, $S(L/2)$ as a function of $\lambda$ for different $\mathbb{Z}_{2}$ electric field values. $\mathbb{Z}_{2}$ LGT results are presented in sub-figures (a)–(d) where the top row shows the entanglement entropy as a function of filling $n$ and $\lambda$, for $h=0$, $h/t=0.4$, $h/t=1$, and $h/t=2$, respectively. The second row of (a)–(d) shows the same results but as a function the chemical potential, instead of filling $n$. The mean-field theory entanglement entropy (in the gauge sector) as a function of filling $n$ and $\lambda$ for $h=0$, $h/t=0.4$, $h/t=1$, and $h/t=2$ are shown in the third row labeled (e)–(h), respectively.
@@ D.2 DMRG simulation of the mean-field model
    We simulate the spin model Eq. (11) by using DMRG. We use $g$ which we calculated in the previous step and choose a $\mathbb{Z}_{2}$ electric field strength $h$. The only missing parameter now is $\mu_{\tau}$. To find the correct $\mu_{\tau}$ which corresponds to the correct filling $n$ we run the following procedure:
@@
      For the given $\mu_{p}$ we calculate the filling by using Eq. (33). If the target filling $n$ is greater than the calculated filling $n(p=1)$ we redefine $\mu_{max}=\mu_{j}$. In contrast if the target filling is smaller than the calculated filling we redefine $\mu_{min}=\mu_{j}$.
@@ D.3 Ground state energy comparison
Figure 16: Central charge results extracted by fitting the entanglement entropy $\tilde{S}(x)$ with Eq. (60) and Eq. (61) for the $\mathbb{Z}_{2}$ LGT Eq. (1) and the mean-field theory Eq. (9). (a) $\mathbb{Z}_{2}$ LGT results for zero $\mathbb{Z}_{2}$ electric field $h=0$. Without discarding fit results with lower quality, we obtain a more pronounced diagonal lines as a function of $n$ for $\lambda\neq 0$, when the $\mathbb{Z}_{2}$ field is zero $h=0$. This agrees with the argument that the entanglement entropy results become more sensitive when the system approaches the criticality. (b) When the $\mathbb{Z}_{2}$ field is non-zero $h/t=1$, the $\mathbb{Z}_{2}$ LGT results show only one line as a function of $n$ for $\lambda\neq 0$. Without discarding lower quality fits, we also see some outliers for $n\geq 0.8$. (c) Mean-field theory for $h=0$, exhibits two vertical lines related to the transition between the topological and trivial states. Note that also here we present data points coming from fits with lower quality, which results in higher maximum value for $c$. (d) The results for the mean-field theory $h/t=1$, show only one transition at $n\approx 0.8$, we note that the maximum value of $c$ is similar to the main text results.
@@ E.1 Entanglement entropy
In the main text we focused on entanglement entropy as a function of filling $n$, $\lambda$, and the $\mathbb{Z}_{2}$ electric field $h$. Here we also present the results for the $\mathbb{Z}_{2}$ LGT when the same data is plotted as a function of the chemical potential $\mu$, see the middle plots in Fig. 14. For the zero electric field $h=0$, a clear qualitative change in the entanglement entropy can be observed at $\mu=\pm 2t$. Such change can be seen for both $\lambda>0$ and $\lambda<0$ albeit with the difference in the absolute magnitude of the entanglement entropy, see also discussion in the main text. The fact that we see a qualitative change at $\mu=\pm 2t$ shows that the simulated model (59) for $h=0$ indeed resembles the behavior of the generic one-dimensional superconducting model (24), where a topological trivial to non trivial transition occurs for $\lambda\neq 0$ at $\mu=\pm 2t$ [28, 35].
Similar results can also be seen for the non-zero values of the $\mathbb{Z}_{2}$ electric field with two differences: the transition at $\mu=-2t$ disappears or at least the change in behavior of the entanglement entropy is smooth, and the transition at $\mu=+2t$ is slightly shifted to a higher value, as a function of $h$, see Fig. 14(b)–(d).

### v2

@@ A.2 Including the $\mathbb{Z}_{2}$ electric field term $h$
$\hat{\tau}^{x}_{\langle j,j+1\rangle}=e^{i\pi\sum_{l<j}\hat{n}_{l}}.$ (27)
[…]
@@ A.3 Mapping between the Kitaev chain and the transverse-field Ising model
$\hat{\mathcal{H}}_{I}=-J\sum_{j}\hat{\sigma}^{x}_{j}\hat{\sigma}^{x}_{j+1}-h_{z}\sum_{j}\hat{\sigma}^{z}_{j},$ (29)
@@
$\hat{\sigma}^{z}_{j}\rightarrow\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle},\quad\hat{\sigma}^{x}_{j}\hat{\sigma}^{x}_{j+1}\rightarrow\hat{\tau}^{z}_{\langle j,j+1\rangle},$ (30)
@@
$\hat{\mathcal{H}}_{I}=-J\sum_{j}\hat{\tau}^{z}_{\langle j,j+1\rangle}-h_{z}\sum_{j}\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}.$ (31)
@@ Appendix B Mapping of the $\mathbb{Z}_{2}$ LGT to the spin-1/2 model
$\hat{n}_{j}=\frac{1}{2}\left(1-\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}\right).$ (33)
@@
$\begin{split}\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{a}^{\dagger}_{j}&=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\frac{1}{2}\left(1+\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}\right),\\ \Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\hat{a}_{j}&=\Bigg(\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}\Bigg)\frac{1}{2}\left(1-\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}\right).\end{split}$ (34)
We once again note that we always consider that our chain starts and ends with a link. The logic of the operator is following: we first take into account that the particle at some site $j$ can be added only if the site $j$ is empty, otherwise the state has to be annihilated. This is achieved by first applying the operator $\frac{1}{2}\left(1+\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}\right)$ which projects to the state with an empty site $j$. After this we apply the product of $\prod_{l<j}\hat{\tau}^{z}_{\langle l,l+1\rangle}$ operators up to the link which attaches to our site from the left. In such way we create a domain wall and thus a particle at that site. The same logic applies for the annihilation operator.
@@
$\hat{\mathcal{H}}_{LGT}^{s}=t\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}-\hat{S}^{z}_{j}\right)-h\sum_{j=1}^{L}2\hat{S}^{x}_{j},$ (35)
@@
$\hat{\mathcal{H}}_{\lambda}^{s}=\lambda\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}+\hat{S}^{z}_{j}\right)$ (36)
and a term which is proportional to the chemical potential $\hat{\mathcal{H}}_{\mu}^{s}=\mu\sum_{j=1}^{L-1}2\hat{S}^{x}_{j}\hat{S}^{x}_{j+1}$. This term is important for the DMRG simulations, since it is used to control the filling in the lattice.
@@ C.1 Derivation of the mean-field models
$\hat{\mathcal{H}}=-t\sum_{j=1}^{L-1}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}_{j}\rangle+\langle\hat{c}^{\dagger}_{j}\hat{c}_{j+1}\rangle\right)\hat{\tau}^{z}_{\langle j,j+1\rangle}+\lambda\sum_{j=1}^{L-1}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}\rangle+\langle\hat{c}_{j}\hat{c}_{j+1}\rangle\right)\hat{\tau}^{z}_{\langle j,j+1\rangle}-h\sum_{j=0}^{L}\hat{\tau}^{x}_{\langle j,j+1\rangle}$
$-t\sum_{j=1}^{L-1}\left(\langle\hat{\tau}^{z}_{\langle j,j+1\rangle}\rangle\hat{c}^{\dagger}_{j+1}\hat{c}_{j}+\text{h.c.}\right)+\lambda\sum_{j=1}^{L-1}\left(\langle\hat{\tau}^{z}_{\langle j,j+1\rangle}\rangle\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}+\text{h.c.}\right),$ (40)
@@
where we defined $t_{c}=t\langle\hat{\tau}^{z}_{\langle j,j+1\rangle}\rangle$ and $\lambda_{c}=\lambda\langle\hat{\tau}^{z}_{\langle j,j+1\rangle}\rangle$, and added the chemical potential term $\mu_{c}$ which enforces the desired filling of the chain.
@@
$\hat{\mathcal{H}}_{\tau}=-g\sum_{j=1}^{L-1}\hat{\tau}^{z}_{\langle j,j+1\rangle}-h\sum_{j=0}^{L}\hat{\tau}^{x}_{\langle j,j+1\rangle}$
$+\mu_{\tau}\sum_{j=1}^{L}\left(\hat{\tau}^{x}_{\langle j-1,j\rangle}\hat{\tau}^{x}_{\langle j,j+1\rangle}-\left(1-2n\right)\right),$ (42)
@@ C.2 Solving the mean-field theory for the charges
$\hat{\mathcal{H}}_{c}=\sum_{k>0}\left(\mu_{c}-2t_{c}\cos(k)\right)\left(\hat{c}^{\dagger}_{k}\hat{c}_{k}+\hat{c}^{\dagger}_{-k}\hat{c}_{-k}\right)$ (45)
@@
$\hat{\mathcal{H}}_{c}=\sum_{k>0}\begin{pmatrix}\hat{c}^{\dagger}_{k}&\hat{c}_{-k}\end{pmatrix}\begin{pmatrix}\epsilon(k)&-2i\lambda_{c}\sin(k)\\ 2i\lambda_{c}\sin(k)&-\epsilon(k)\end{pmatrix}\begin{pmatrix}\hat{c}_{k}\\ \hat{c}^{\dagger}_{-k}\end{pmatrix}$
@@
$\hat{\mathcal{H}}_{c}-\sum_{k>0}\epsilon(k)+\mu_{c}nL=\sum_{k>0}\begin{pmatrix}\hat{c}^{\dagger}_{k}&\hat{c}_{-k}\end{pmatrix}\bm{H}(k)\begin{pmatrix}\hat{c}_{k}\\ \hat{c}^{\dagger}_{-k}\end{pmatrix}\\ =\sum_{k>0}\begin{pmatrix}\hat{b}^{\dagger}_{k}&\hat{b}_{-k}\end{pmatrix}\bm{\Lambda}(k)\begin{pmatrix}\hat{b}_{k}\\ \hat{b}^{\dagger}_{-k}\end{pmatrix}$, (49)
@@
$\hat{\mathcal{H}}_{c}=\sum_{k}\Lambda_{+}\hat{b}^{\dagger}_{k}\hat{b}_{k}-\sum_{k>0}\Lambda_{+}+\sum_{k>0}\epsilon(k)-\mu_{c}nL.$ (51)
@@ C.3 Solving the mean-field theory for the gauge fields
$E_{0}=-t_{c}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}_{j}\rangle+\langle\hat{c}^{\dagger}_{j}\hat{c}_{j+1}\rangle\right)\\ +\lambda_{c}\left(\langle\hat{c}^{\dagger}_{j+1}\hat{c}^{\dagger}_{j}\rangle+\langle\hat{c}_{j}\hat{c}_{j+1}\rangle\right)=-\langle\hat{\tau}^{z}_{\langle j,j+1\rangle}\rangle g.$ (56)
@@ D.1 DMRG simulation of the $\mathbb{Z}_{2}$ LGT
$\hat{\mathcal{H}}_{\rm DMRG}=t\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}-\hat{S}^{z}_{j}\right)\\ +\lambda\sum_{j=2}^{L-1}\left(4\hat{S}^{x}_{j-1}\hat{S}^{x}_{j+1}\hat{S}^{z}_{j}+\hat{S}^{z}_{j}\right)\\ -h\sum_{j=1}^{L}2\hat{S}^{x}_{j}+\mu\sum_{j=1}^{L-1}2\hat{S}^{x}_{j}\hat{S}^{x}_{j+1},$ (59)
@@
Figure 14: Entanglement entropy, $S(L/2)$ as a function of $\lambda$ for different $\mathbb{Z}_{2}$ electric field term values. $\mathbb{Z}_{2}$ LGT results are presented in sub-figures (a)–(d) where the top row shows the entanglement entropy as a function of filling $n$ and $\lambda$, for $h=0$, $h/t=0.4$, $h/t=1$, and $h/t=2$, respectively. The second row of (a)–(d) shows the same results but as a function the chemical potential, instead of filling $n$. The mean-field theory entanglement entropy (in the gauge sector) as a function of filling $n$ and $\lambda$ for $h=0$, $h/t=0.4$, $h/t=1$, and $h/t=2$ are shown in the third row labeled (e)–(h), respectively.
@@ D.2 DMRG simulation of the mean-field model
    We simulate the spin model Eq. (11) by using DMRG. We use $g$ which we calculated in the previous step and choose a $\mathbb{Z}_{2}$ electric field term strength $h$. The only missing parameter now is $\mu_{\tau}$. To find the correct $\mu_{\tau}$ which corresponds to the correct filling $n$ we run the following procedure:
@@
      For the given $\mu_{p}$ we calculate the filling by using Eq. (33). If the target filling $n$ is greater than the calculated filling $n(p=1)$ we redefine $\mu_{min}=\mu_{j}$. In contrast if the target filling is smaller than the calculated filling we redefine $\mu_{max}=\mu_{j}$.
@@ D.3 Ground state energy comparison
Figure 16: Central charge results extracted by fitting the entanglement entropy $\tilde{S}(x)$ with Eq. (60) and Eq. (61) for the $\mathbb{Z}_{2}$ LGT Eq. (1) and the mean-field theory Eq. (9). (a) $\mathbb{Z}_{2}$ LGT results for zero $\mathbb{Z}_{2}$ electric field term $h=0$. Without discarding fit results with lower quality, we obtain a more pronounced diagonal lines as a function of $n$ for $\lambda\neq 0$, when the $\mathbb{Z}_{2}$ field is zero $h=0$. This agrees with the argument that the entanglement entropy results become more sensitive when the system approaches the criticality. (b) When the $\mathbb{Z}_{2}$ field is non-zero $h/t=1$, the $\mathbb{Z}_{2}$ LGT results show only one line as a function of $n$ for $\lambda\neq 0$. Without discarding lower quality fits, we also see some outliers for $n\geq 0.8$. (c) Mean-field theory for $h=0$, exhibits two vertical lines related to the transition between the topological and trivial states. Note that also here we present data points coming from fits with lower quality, which results in higher maximum value for $c$. (d) The results for the mean-field theory $h/t=1$, show only one transition at $n\approx 0.8$, we note that the maximum value of $c$ is similar to the main text results.
@@ E.1 Entanglement entropy
In the main text we focused on entanglement entropy as a function of filling $n$, $\lambda$, and the $\mathbb{Z}_{2}$ electric field term $h$. Here we also present the results for the $\mathbb{Z}_{2}$ LGT when the same data is plotted as a function of the chemical potential $\mu$, see the middle plots in Fig. 14. For the zero electric field term $h=0$, a clear qualitative change in the entanglement entropy can be observed at $\mu=\pm 2t$. Such change can be seen for both $\lambda>0$ and $\lambda<0$ albeit with the difference in the absolute magnitude of the entanglement entropy, see also discussion in the main text. The fact that we see a qualitative change at $\mu=\pm 2t$ shows that the simulated model (59) for $h=0$ indeed resembles the behavior of the generic one-dimensional superconducting model (24), where a topological trivial to non trivial transition occurs for $\lambda\neq 0$ at $\mu=\pm 2t$ [28, 35].
Similar results can also be seen for the non-zero values of the $\mathbb{Z}_{2}$ electric field term with two differences: the transition at $\mu=-2t$ disappears or at least the change in behavior of the entanglement entropy is smooth, and the transition at $\mu=+2t$ is slightly shifted to a higher value, as a function of $h$, see Fig. 14(b)–(d).